Fermi National Laboratory
 

Operation Summary for February 18, 2011 - February 21, 2011

 

The Integrated Luminosity for the period from 2/14/11 to 2/21/11 was 33.64 pb-1 (inverse picobarns). NuMI reported receiving 7.42E18 protons on target during this same period.

Friday Feb 18

The day shift began with Operations monitoring the Tevatron (TEV) store 8510 (18.88E30) and the Recycler stash (151.6E10), with the Antiproton Source (Pbar) stacking (58.2E10), with the Pelletron work continuing, with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported some beam position problems with store 8510 at sector D0. TEV personnel reported a vacuum leak on the D17 collimators. Operations terminated the store and prepared the TEV for access to begin repairs. Operations reported the completion of the Pelletron work; Electron Cooling (ECool) personnel began filling the Pelletron tank with the insulating gas (SF6). When that is finished, the gas has to have its water removed

During the evening shift, Vacuum personnel reported completing their TEV vacuum work at D17. Operators put the TEV into HEP shot setup at the end of shift. The Pelletron gas drying continued.

Saturday Feb 19

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (62.5E10), with Pbar stacking (238.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 8514 with an average initial luminosity (AIL) of 109.7E30. Operations reported that the Pelletron turn on continued until 7 AM, when a problem developed requiring the removal of the SF6 for another Pelletron tank access.

During the day shift, Operations terminated store 8514 early in the shift. Operations established store 8515 at 10 AM with an AIL of 109.6E30.

During the evening shift, Operations terminated store 8515 at 6 PM. Operations established store 8516 at 7:15 PM with an AIL of 103.4E30. Operations reported that the ECool personnel had completed their Pelletron repairs and began filling the tank with the SF6 again.

Sunday Feb 20

The midnight shift began with Operations monitoring store 8516 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8516 at 3 AM. The Pelletron came online at 4 AM. Operations established store 8517 at 4:15 AM with an AIL of 110.5E30.

During the day shift, Operations terminated store 8517 at noon. Operations established store 8518 at 1:20 PM with an AIL of 280.5E30. Operators reported a Linac quadrupole magnet power supply had failed, which interrupted beam for about three hours.

During the evening shift, Operations monitored the Accelerator Complex.

Monday Feb 21

The midnight shift began with Operations monitoring store 8518 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that stacking was interrupted for about 45 minutes due to a Pbar LCW flow sensor.

The Plans for Monday

The plans for today are to establish a new store this morning. After the store is established, Pbar personnel will access and begin working on their dump LCW leaks; they estimated that this job would take about 18 hours.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/21/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory